OpenStudy (anonymous):

Explain: 3(y+3)^2=27

OpenStudy (anonymous):

Firstly divide both sides by 3 to get (y+3)^2=9 Then take the square root to get y+3=3 and y+3=-3 Solving both equation we get that y equals 0 or -6

OpenStudy (anonymous):

Thank you.